OpenStudy (anonymous):

A die is rolled 7 times. What is the probability of rolling a perfet square each time?

OpenStudy (anonymous):

i guess the only perfect square on one die is 1 and 4, so the probability you roll a perfect square one time is \(\frac{2}{6}=\frac{1}{3}\) therefore the probability you do it 7 times in a row is \[\left(\frac{1}{3}\right)^7\]

OpenStudy (anonymous):

oooh thank you. I didn't understand what a perfect square was